Say True or False with proof



Q1. If 𝑓(π‘₯) = 2|π‘₯ βˆ’ 1| and 𝑔(π‘₯) = 3π‘₯ βˆ’ 10,


then π‘“π‘œπ‘”(1) = 12.


Q2. (√2, 1,1/2)∈ 𝑸 Γ— 𝒁 Γ— 𝑹.


Q3. The domain of the function f defined by 𝑓(π‘₯)


=√{(3-x)/(x-2)} is R-{2}.

1
Expert's answer
2022-04-16T05:39:33-0400

Q1 false

"f\circ g= 2|3x-10-1|=2|3x-11|"

"f\circ g(1)=2|3\cdot 1-11|=16"


Q2 false

"\sqrt 2\notin Q,1\isin Z,1/2\isin R"


Q3 false

"f(x)=\sqrt{(3-x)/(x-2)}"

domain: "x\isin (2,3]"